sum of two numbers is 42 and their difference is 30. find each of the numbers.
Answer by ikdeep(189)

Let the first number = x
Let the second number = y
According first condition i.e sum of two numbers is 42 ...we get
x+y = 42............(1)
According second condition i.e difference of two numbers is 30 ...we get
x-y = 30.............(2)
on adding (1)and (2)..we get...

x+y = 42
x-y = 30
--------
2x = 72 (+y and -y would be eliminated)
divide both side by 2 we get...
x = 72/2
x = 36
,,,to get the value of y, substitute the value of x either in (1) or in (2)...but here i took (1)....
x+y = 42............(1)
36 + y = 42
y = 42 - 36
y = 6
so the values are x = 36 and y = 6......
To verify your answer you can put the value of x and y in either (1) or(2)
or in both and you will always get LHS = RHS
please feel free to revert back for any further queries.